OpenStudy (anonymous):

5(x - 5) = 4(x - 1) + 3x.

OpenStudy (anonymous):

5x - 25 = 4x - 4 + 3x 5x = 21 + 7x -2x = 21 x = 10 1/2 or - 21/2